for a random sample of 36 items and a sample mean of 211,compute a 95% confidence interval for if population standard deviation is 23



1
Expert's answer
2021-12-20T13:33:51-0500

The z value is; 1.96


The interval will be calculated by the following formula;


"mean \pm \frac{z×sd}{\sqrt{N}}"


"211 \pm \frac{1.96×23}{\sqrt{36}}"


Evaluating we get an interval of:


203.49 to 218.51